OpenStudy (anonymous):

Choose one of the factors of 500x3 + 108y18 (500x to the 3rd + 108y to the 18th) A. 25x2 + 15xy6 + 9y2 B. 5x + 3y6 C. All of the above D. 6

OpenStudy (anonymous):

A&B

OpenStudy (anonymous):

4(5x + 3y6)(25x2 + 15xy6 + 9y2)

OpenStudy (anonymous):

500x^3 + 108y^18 =4(125x^3+27y^18) =4(5^3*x^3+3^3*(y^6)^3) =4[(5x)^3+(3y^6)^3] =4(5x+3y^6)(25x^2-18xy^6+9y^12) answer is B

OpenStudy (anonymous):

\[500x{}^{\wedge}3 + 108y{}^{\wedge}18=4 \left(5 x+3 y^6\right) \left(25 x^2-15 x y^6+9 y^{12}\right) \]B.